# Guess your neighbor’s input: Quantum advantage in Feige’s game

### v1

@@ Abstract.
In this article, we study a nonlocal game with two questions and three answers per player, which was first considered by Feige in 1991, and show that there is quantum advantage in this game. We prove that the game is a robust self-test for the 33-dimensional maximally entangled state. Furthermore, we show that the game can be seen as the ”or” of two games that each do not have quantum advantage. Lastly, we investigate the behavior of the game with respect to parallel repetition in the classical, quantum and non-signalling case and obtain perfect parallel repetition of the non-signalling value if Feige’s game is repeated an even amount of times.
@@ 1. Introduction
The same relations hold for Bob’s operators. Note that the relations look similar to the relations obtained from mutually unbiased measurements TFR+ (21), where for the latter it holds $\nu=\frac{1}{d}I$. In case of $P^{x}_{a}=\mathinner{|\phi^{x}_{a}\rangle}\mathinner{\langle\phi_{a}^{x}|}$ being rank 11 projections, the transition matrix from the basis $\{\mathinner{|\phi^{x}_{a}\rangle}\}$ to the basis $\{\mathinner{|\phi^{1-x}_{a}\rangle}\}$ is given by a unitary matrix $H$ with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ (which is a Hadamard matrix in the case of $\nu=\frac{1}{d}I$). Similar to some Hadamard matrices being unique up to equivalence Haa (97), we show that the same is true for unitaries with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ for the specific choice of $\nu$ from above. More generally, we give simple criterion for when such a unitary is unique up to a certain kind of equivalence. Using this, and the operator-algebraic framework for self-testing developed in Zha (24), we obtain that Feige’s game is a self-test. Another ingredient we need in our proof is that in an optimal strategy, the actions of Bob’s projective measurements on the state are determined by actions of (self-adjoint polynomials in) Alice’s projective measurements.
@@
Does the quantum value exceed the classical value for the 33-fold parallel repetition of Feige’s game?
We are not aware of a game with quantum advantage, for which the classical and quantum value agree for the 22-fold parallel repetition, but then they separate again for $n=3$. As we can see in Table 1 this happens for the classical and the non-signalling value of Feige’s game. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Therefore, there is no natural candidate of a quantum strategy that yields quantum advantage for the 33-fold parallel repetition. The $\frac{5}{16}$ optimal classical strategy is also interesting: It performs worse on the first two rounds of the game in comparison to the optimal strategy for the 22-fold parallel repetition, but is constructed such that if the players win the first two rounds, they win the third round automatically.
@@ 2.1. Strategies for nonlocal games
    probability distributions $\{p_{\omega}(a|x):x\in X,\omega\in\Omega\}$ with outcomes in AA and $\{q_{\omega}(b|y):y\in Y,\omega\in\Omega\}$ with outcomes in $B$.
@@
The proof of our main theorem follows an operator-algebraic framework for self-testing developed inZha (24). For finite sets $X$ and AA, we use $\mathscr{A}_{PVM}^{X,A}$ to denote the universal $C^{*}$-algebra generated by $\{p^{x}_{a}:a\in A,x\in X\}$, subject to the relations
@@ 3. Feige’s game
    Once a player answers a bit given a question (e.g. Alice answers 0 when receiving 0), the game will always be lost if the other player receives a question different from this bit (the referee sends 11 to Bob), which happens halve of the time.
@@ 3.2. Method for obtaining an upper bounds and algebraic relations
$3.$ The non-signalling value of Feige’s game
@@ 4. Feige’s game is a self-test
where $\nu$ is the symmetric stochastic matrix defined in $4.$ In fact, for any $n\times n$ symmetric stochastic matrix $\nu$, we can define $\mathcal{A}_{\nu}$ analogously, where generators $\{E_{i}\}_{i=1}^{n}$ and $\{F_{i}\}_{i=1}^{n}$ are PVMs satisfying $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$ and $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$. In Appendix B, we establish the representation theory for those $\mathcal{A}_{\nu}$’s. Interestingly, we show that representations for $\mathcal{A}_{\nu}$ can be seen as a generalization of mutually unbiased measurements (MUMs, see e.g. TFR+ (21)). We generalize the results in TFR+ (21) and give a characterization of irreducible representations of $\mathcal{A}_{\nu}$. In particular, in B.9, we show that for $2\times 2$ and $3\times 3$ symmetric stochastic matrices $\nu$, every $\mathcal{A}_{\nu}$ has a unique irreducible representation up to unitary equivalence and relabeling generators. For the matrix $\nu$ defined in 4.4, we further show that any allowed relabeling still results in a unitarily equivalent representation.
@@
Let $(\Gamma,\mathcal{R})$ be the $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$ defined in $4.$ By 4.5, the associated game algebra $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is a full matrix algebra. It follows from 4.3 that $G_{F}$ is a robust self-test.
@@ 6. Feige’s game and parallel repetition
In this section, we look at parallel repetition of Feige’s game. From Table 1, we see that Feige’s game has interesting behaviour with respect to parallel repetition. In the table, the classical, quantum and non-signalling values agree in the case of an even number of repetitions, but this is not the case for odd numbers. As we saw previously, all values are different for the original game. Furthermore, we get that at least the classical and non-signalling value are different for $n=3$, despite this not being the case for $n=2$. We do not know if there is quantum advantage for $n=3$. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
@@ 6.1. Classical strategies for the $2m$-fold parallel repetition
respectively. We see that the players win if and only if $x_{1}=y_{2}$, which happens with probability $\frac{1}{2}$. This may be surprising, since the classical winning probability of the 22-fold parallel repetition does not decrease when compared with the original game $G_{F}$.
@@ 6.2. The non-signalling upper bound for $2m$-fold parallel repetition
For this, we first show the following lemma. We define $\mathbf{x}|_{I}\in\{0,1\}^{|I|}$ to be the restriction of $\mathbf{x}$ to a subset $I\subseteq[n]$. Furthermore, we write $|I|$ for the cardinality of a set II and denote the complement by $I^{c}=[n]\setminus I$.
@@
because of $p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ and the non-signalling property of pp. Since $b_{I}(\mathbf{v})=b_{I}(\mathbf{w})$ if and only if $\mathbf{v}|_{I}=\mathbf{w}|_{I}$, we see that the right hand side of Equation 15 is the same for each $\mathbf{x}|_{I^{c}}\in J$. This yields the assertion. ∎
@@
Since $n=2m$, both sets II and $I^{c}$ have $m$ elements. Therefore, we can choose pairs $(i_{k},j_{k})$ for $1\leq k\leq m$ with $i_{k}\in I,j_{k}\in I^{c}$ such that $\{i_{k},j_{k}|1\leq k\leq m\}=[n]$. It holds
@@
for any nonsignalling strategy pp. From 6.3, we obtain for $|I|>m$ that
@@
6.3. Values for the 33-fold parallel repetition
Now, we will look at the case of 33-fold parallel repetition. We will start with describing a classical strategy with winning probability $\frac{5}{16}$. We note that this strategy performs better that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Alice’s and Bob’s strategies are of the form $\left(a_{1},a_{2},a_{3}\right)$, $\left(b_{1},b_{2},b_{3}\right)$ where
@@ 6.3. Values for the 33-fold parallel repetition
If $x_{1}=y_{2}\wedge y_{3}=0$, the players win round 33 if and only if $y_{2}=x_{3}$. But this is true because $y_{2}=x_{1}\vee x_{3}=x_{3}$ for $x_{1}=0$. If $x_{1}=y_{2}\wedge y_{3}=1$, the players win if and only if $y_{3}=1$, which is true since $y_{2}\wedge y_{3}=1$ implies $y_{3}=1$. This shows $\Pr[W_{2}|W_{1},W_{3}]=1$ and by Equation 20, the strategy has winning probability $\frac{5}{16}$. By exhaustive search through all deterministic strategies using a computer, we have confirmed that this strategy is indeed optimal.
While we see that the previous classical strategy performs better than the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the 22-fold parallel repetition, there is no non-signalling strategy that outperforms the one where we use optimal non-signalling strategy for one game and the optimal classical strategy for the 22-fold parallel repetition. Therefore, the non-signalling value of the 33-fold parallel repetition is $\frac{1}{3}$. The following is an optimal strategy achieving this winning probability.

### v2

@@ Abstract.
In this article, we study a nonlocal game with two questions and three answers per player, which was first considered by Feige in 1991, and show that there is quantum advantage in this game. We prove that the game is a robust self-test for the $3$-dimensional maximally entangled state. Furthermore, we show that the game can be seen as the ”or” of two games that each do not have quantum advantage. Lastly, we investigate the behavior of the game with respect to parallel repetition in the classical, quantum and non-signalling case and obtain perfect parallel repetition of the non-signalling value if Feige’s game is repeated an even amount of times.
@@ 1. Introduction
The same relations hold for Bob’s operators. Note that the relations look similar to the relations obtained from mutually unbiased measurements TFR+ (21), where for the latter it holds $\nu=\frac{1}{d}I$. In case of $P^{x}_{a}=\mathinner{|\phi^{x}_{a}\rangle}\mathinner{\langle\phi_{a}^{x}|}$ being rank $1$ projections, the transition matrix from the basis $\{\mathinner{|\phi^{x}_{a}\rangle}\}$ to the basis $\{\mathinner{|\phi^{1-x}_{a}\rangle}\}$ is given by a unitary matrix $H$ with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ (which is a Hadamard matrix in the case of $\nu=\frac{1}{d}I$). Similar to some Hadamard matrices being unique up to equivalence Haa (97), we show that the same is true for unitaries with $|h_{aa^{\prime}}|=\sqrt{\nu_{aa^{\prime}}}$ for the specific choice of $\nu$ from above. More generally, we give simple criterion for when such a unitary is unique up to a certain kind of equivalence. Using this, and the operator-algebraic framework for self-testing developed in Zha (24), we obtain that Feige’s game is a self-test. Another ingredient we need in our proof is that in an optimal strategy, the actions of Bob’s projective measurements on the state are determined by actions of (self-adjoint polynomials in) Alice’s projective measurements.
@@
Does the quantum value exceed the classical value for the $3$-fold parallel repetition of Feige’s game?
We are not aware of a game with quantum advantage, for which the classical and quantum value agree for the $2$-fold parallel repetition, but then they separate again for $n=3$. As we can see in Table 1 this happens for the classical and the non-signalling value of Feige’s game. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Therefore, there is no natural candidate of a quantum strategy that yields quantum advantage for the $3$-fold parallel repetition. The $\frac{5}{16}$ optimal classical strategy is also interesting: It performs worse on the first two rounds of the game in comparison to the optimal strategy for the $2$-fold parallel repetition, but is constructed such that if the players win the first two rounds, they win the third round automatically.
@@ 2.1. Strategies for nonlocal games
    probability distributions $\{p_{\omega}(a|x):x\in X,\omega\in\Omega\}$ with outcomes in $A$ and $\{q_{\omega}(b|y):y\in Y,\omega\in\Omega\}$ with outcomes in $B$.
@@
The proof of our main theorem follows an operator-algebraic framework for self-testing developed inZha (24). For finite sets $X$ and $A$, we use $\mathscr{A}_{PVM}^{X,A}$ to denote the universal $C^{*}$-algebra generated by $\{p^{x}_{a}:a\in A,x\in X\}$, subject to the relations
@@ 3. Feige’s game
    Once a player answers a bit given a question (e.g. Alice answers 0 when receiving 0), the game will always be lost if the other player receives a question different from this bit (the referee sends $1$ to Bob), which happens halve of the time.
@@ 3.2. Method for obtaining an upper bounds and algebraic relations
3.3. The non-signalling value of Feige’s game
@@ 4. Feige’s game is a self-test
where $\nu$ is the symmetric stochastic matrix defined in 4.4. In fact, for any $n\times n$ symmetric stochastic matrix $\nu$, we can define $\mathcal{A}_{\nu}$ analogously, where generators $\{E_{i}\}_{i=1}^{n}$ and $\{F_{i}\}_{i=1}^{n}$ are PVMs satisfying $E_{i}F_{j}E_{i}=\nu_{ij}E_{i}$ and $F_{i}E_{j}F_{i}=\nu_{ij}F_{i}$. In Appendix B, we establish the representation theory for those $\mathcal{A}_{\nu}$’s. Interestingly, we show that representations for $\mathcal{A}_{\nu}$ can be seen as a generalization of mutually unbiased measurements (MUMs, see e.g. TFR+ (21)). We generalize the results in TFR+ (21) and give a characterization of irreducible representations of $\mathcal{A}_{\nu}$. In particular, in B.9, we show that for $2\times 2$ and $3\times 3$ symmetric stochastic matrices $\nu$, every $\mathcal{A}_{\nu}$ has a unique irreducible representation up to unitary equivalence and relabeling generators. For the matrix $\nu$ defined in 4.4, we further show that any allowed relabeling still results in a unitarily equivalent representation.
@@
Let $(\Gamma,\mathcal{R})$ be the $O(\sqrt{\epsilon})$-robust determining pair of $G_{F}$ defined in 4.4. By 4.5, the associated game algebra $C^{*}(G)=\mathscr{A}_{PVM}^{\{0,1\},\{0,1,\bot\}}/\langle\mathcal{R}\rangle$ is a full matrix algebra. It follows from 4.3 that $G_{F}$ is a robust self-test.
@@ 6. Feige’s game and parallel repetition
In this section, we look at parallel repetition of Feige’s game. From Table 1, we see that Feige’s game has interesting behaviour with respect to parallel repetition. In the table, the classical, quantum and non-signalling values agree in the case of an even number of repetitions, but this is not the case for odd numbers. As we saw previously, all values are different for the original game. Furthermore, we get that at least the classical and non-signalling value are different for $n=3$, despite this not being the case for $n=2$. We do not know if there is quantum advantage for $n=3$. Note that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition performs worse than the optimal classical strategy for $n=3$, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$.
@@ 6.1. Classical strategies for the $2m$-fold parallel repetition
respectively. We see that the players win if and only if $x_{1}=y_{2}$, which happens with probability $\frac{1}{2}$. This may be surprising, since the classical winning probability of the $2$-fold parallel repetition does not decrease when compared with the original game $G_{F}$.
@@ 6.2. The non-signalling upper bound for $2m$-fold parallel repetition
For this, we first show the following lemma. We define $\mathbf{x}|_{I}\in\{0,1\}^{|I|}$ to be the restriction of $\mathbf{x}$ to a subset $I\subseteq[n]$. Furthermore, we write $|I|$ for the cardinality of a set $I$ and denote the complement by $I^{c}=[n]\setminus I$.
@@
because of $p(a_{I}(\mathbf{y}),b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})\leq\sum_{\mathbf{a}}p(\mathbf{a},b_{I}(\mathbf{x})|\mathbf{x},\mathbf{y})$ and the non-signalling property of $p$. Since $b_{I}(\mathbf{v})=b_{I}(\mathbf{w})$ if and only if $\mathbf{v}|_{I}=\mathbf{w}|_{I}$, we see that the right hand side of Equation 15 is the same for each $\mathbf{x}|_{I^{c}}\in J$. This yields the assertion. ∎
@@
Since $n=2m$, both sets $I$ and $I^{c}$ have $m$ elements. Therefore, we can choose pairs $(i_{k},j_{k})$ for $1\leq k\leq m$ with $i_{k}\in I,j_{k}\in I^{c}$ such that $\{i_{k},j_{k}|1\leq k\leq m\}=[n]$. It holds
@@
for any nonsignalling strategy $p$. From 6.3, we obtain for $|I|>m$ that
@@
6.3. Values for the $3$-fold parallel repetition
Now, we will look at the case of $3$-fold parallel repetition. We will start with describing a classical strategy with winning probability $\frac{5}{16}$. We note that this strategy performs better that the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition, since $\frac{9}{16}\cdot\frac{1}{2}=\frac{9}{32}<\frac{5}{16}$. Alice’s and Bob’s strategies are of the form $\left(a_{1},a_{2},a_{3}\right)$, $\left(b_{1},b_{2},b_{3}\right)$ where
@@ 6.3. Values for the $3$-fold parallel repetition
If $x_{1}=y_{2}\wedge y_{3}=0$, the players win round $3$ if and only if $y_{2}=x_{3}$. But this is true because $y_{2}=x_{1}\vee x_{3}=x_{3}$ for $x_{1}=0$. If $x_{1}=y_{2}\wedge y_{3}=1$, the players win if and only if $y_{3}=1$, which is true since $y_{2}\wedge y_{3}=1$ implies $y_{3}=1$. This shows $\Pr[W_{2}|W_{1},W_{3}]=1$ and by Equation 20, the strategy has winning probability $\frac{5}{16}$. By exhaustive search through all deterministic strategies using a computer, we have confirmed that this strategy is indeed optimal.
While we see that the previous classical strategy performs better than the quantum strategy consisting of the optimal quantum strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition, there is no non-signalling strategy that outperforms the one where we use optimal non-signalling strategy for one game and the optimal classical strategy for the $2$-fold parallel repetition. Therefore, the non-signalling value of the $3$-fold parallel repetition is $\frac{1}{3}$. The following is an optimal strategy achieving this winning probability.